# Think, Act, Learn: A Framework for Autonomous Robotic Agents using Closed-Loop Large Language Models

### v1

@@ II-E Benchmarks and Evaluation
Rigorous evaluation is the bedrock of scientific progress. The agent community has made significant strides in developing challenging benchmarks and realistic evaluation environments. For web agents, WebArena [3] provides a realistic and reproducible environment, while the Mind2Web dataset [2] offers a diverse set of tasks. The evaluation has since expanded to mobile [10, 11] and desktop environments [12]. To assess the general capabilities of foundation models as assistants, comprehensive benchmarks like GAIA [40] and AgentBench [39] have been proposed. Other benchmarks target specific skills, such as instruction-following with vision [30], mathematical reasoning from visual data [38], or recognizing micro-actions [29]. The insights from these evaluations are crucial; for example, empirical studies have systematically analyzed the failure modes of current agents on real-world tasks [6]. Furthermore, the application of LLM-based agents to software engineering tasks like UI testing is an active area of investigation [46], as is the potential for automating UI-to-code generation [52]. Our work contributes to this ecosystem by evaluating our proposed agent on a diverse set of tasks and comparing it against relevant state-of-the-art methods, following the principles laid out in these foundational evaluation efforts. We also draw inspiration from work that transfers knowledge from large-scale, unstructured data like internet videos to embodied control [53], a paradigm that holds promise for teaching agents from the vast repository of software tutorials online. Finally, incorporating human-like interaction patterns, such as those informed by gaze data [59], represents a promising direction for making agents more intuitive and efficient.
@@ III-C Perception Module: Structured UI Representation
The module is built upon a powerful pre-trained Vision-Language Model (VLM), inspired by recent advancements in multimodal understanding [18, 19]. The input to the VLM is a prompt that combines the task goal $g$ and the screenshot $I_{t}$. The model is tasked with identifying all interactable elements on the screen. The output is a structured list of UI elements, $U_{t}$, which forms the agent’s symbolic understanding of the current interface state. Each element $e_{i}$ in this list is a tuple containing its essential properties:
@@ III-D Decision Module: Action Generation
The architecture of this module is a Transformer decoder [25]. The input to the Transformer is a sequence of tokens constructed by serializing the information from the Perception Module and the task history. Specifically, the structured representation $U_{t}$ is flattened into a sequence of strings, where each element $e_{i}$ is described as, for example, ‘”[ID1] Button ’Login’ at (120, 300, 200, 340)”‘. This sequence is concatenated with the tokenized goal $g$ and a summary of the recent history $h_{t}$.
@@
[…]
@@
[…]
@@ III-E2 Phase 2: Offline Reinforcement Learning from Self-Exploration
With the labeled offline dataset $\mathcal{D}_{O}$, we use an offline RL algorithm to update the policy. We specifically choose a conservative, actor-critic algorithm like Implicit Q-Learning (IQL) [35], which is well-suited for learning from static datasets without querying the environment. IQL avoids explicitly maximizing the Q-function, which can be unstable with out-of-distribution actions, and instead learns by expecting the Q-values for actions in the dataset to be high. It involves learning a state-value function $V_{\psi}$ and a Q-function $Q_{\phi}$ using an expectile regression loss. The policy $\pi_{\theta}$ is then updated to maximize the exponentiated advantage term, pushing the policy towards actions with higher Q-values relative to the average value of the state. The policy update loss is given by:
@@ IV-A1 Datasets and Environments
  • Mind2Web [2]: This dataset provides a large collection of tasks across various websites, focusing on generalist web agents. We use a subset of its tasks for both training (behavioral cloning) and zero-shot evaluation to test generalization to unseen websites within the same domain.
  • WebArena [3]: This benchmark provides a realistic and reproducible web environment with tasks that often require complex reasoning and interaction sequences (e.g., online shopping, content management). We use WebArena primarily for evaluation to test the agent’s ability to handle dynamic and complex web applications.
@@
To demonstrate the platform-agnostic nature of our pixel-based approach, we also conduct a qualitative evaluation on tasks from the Android-in-the-Wild dataset [11], showcasing the agent’s ability to operate on a mobile GUI without any architectural changes.
@@ IV-A2 Implementation Details
We use a pre-trained Vision-Language Model as the backbone. Specifically, we initialize our model with weights from CogAgent-VQA-hf [19], which is optimized for GUI understanding and visual question answering. We further fine-tune this model during the behavioral cloning phase to specialize it for our structured UI representation extraction task. The input image resolution is set to 1120x1120 pixels.
@@ V-A Main Quantitative Results
We evaluated GUI-Learner against a suite of strong baselines to demonstrate its effectiveness. The baselines include: (1) Behavioral Cloning (BC), which represents our model trained only with the supervised learning phase (Phase 1); (2) GPT-4V (Zero-shot), where we use the official OpenAI API with a carefully crafted prompt based on the ReAct framework [26] to directly generate actions from screenshots; and (3) Pix2Act [33], a representative earlier vision-based agent that maps pixels directly to actions. The main results, measured in Task Success Rate (SR) and Action Efficiency (AE), are summarized in Table I.
@@
GPT-4V (Zero-shot) [18] 41.3 0.68 35.7 0.61
Pix2Act [33] 32.5 0.55 28.1 0.52
@@
As shown in Table I, GUI-Learner achieves state-of-the-art performance across all environments and metrics. On the web environments, combining tasks from Mind2Web [2] and WebArena [3], our agent achieves a success rate of 82.4%, a significant improvement of 16.6 percentage points over the strong Behavioral Cloning baseline. This underscores the critical role of the offline reinforcement learning phase. While BC provides a strong initial policy, it struggles with the problem of covariate shift; when the agent makes a small mistake and enters a state not seen in the expert data, it often cannot recover. Our RL phase, by learning from a diverse dataset of self-generated trajectories including failures, equips the agent with a more robust policy capable of navigating out of these unfamiliar states. This ability to self-correct is a hallmark of more advanced agentic frameworks like Reflexion [36], and our results confirm its importance in the UI domain.
@@
The results on the desktop environments mirror the trends observed on the web. GUI-Learner maintains a substantial performance gap over the baselines, demonstrating that our approach is platform-agnostic. This is a key advantage of our vision-based methodology, which, unlike approaches reliant on web-specific structures like HTML, can seamlessly transition between web, desktop [12], and even mobile environments [10]. The challenges in desktop automation are distinct, often involving interaction between multiple applications and the operating system shell, a direction explored in recent pilot studies [13]. Our strong performance here suggests that GUI-Learner is a significant step towards a truly universal computer control agent [14].
@@ V-B Ablation Studies
The most significant finding from our ablation study is the impact of the offline reinforcement learning phase. The ”w/o Offline RL Phase” variant, which is equivalent to our BC baseline, sees its success rate drop by 17.4 points. This confirms our hypothesis that while imitation learning provides an excellent starting point, it is insufficient for achieving high robustness. The RL phase allows the agent to learn from its own mistakes, effectively expanding its knowledge beyond the limited scope of the expert demonstration dataset $\mathcal{D}_{E}$. This self-improvement loop, where an agent learns from its own experience, is a central theme in modern agent research, whether it’s through explicit reflection [36] or interaction with a human in the loop [37]. Our results provide strong evidence for its efficacy in the context of GUI automation.
@@
In the ”w/o Structured Representation” variant, we replaced our two-module design with a single end-to-end model that maps raw pixels to action coordinates, similar in spirit to older methods [33] but built on a modern VLM backbone. The performance plummets to a 45.5% success rate. This drastic drop highlights the critical role of the Perception Module’s abstraction. By converting pixels into a symbolic, object-centric representation, we provide the Decision Module with a much cleaner and more structured input space. This abstraction allows the policy to be invariant to trivial visual changes (e.g., color themes, font styles, minor layout shifts) and to focus on the functional semantics of the interface. This mirrors findings in related domains, where structured representations like graph inference [22] or program synthesis [54] have been shown to improve reasoning and generalization.
@@
The ”w/o Expert Pre-training” variant represents an attempt to train the agent using pure reinforcement learning from scratch. The result is a near-complete failure, with a success rate of only 12.8%. This is expected in a domain with such a vast state space and sparse rewards. Without the initial guidance provided by behavioral cloning, the agent’s random exploration is highly unlikely to stumble upon a successful trajectory for any non-trivial task. This finding underscores the sample inefficiency of pure RL in complex interactive domains and validates our choice of a hybrid learning strategy. Bootstrapping from expert data is a common and effective technique, not just in UI automation [32] but also in robotics, where learning from internet-scale video data provides a similar warm start [53].
@@ V-C Generalization to Unseen Applications
GUI-Learner achieves a generalization score of 68.9%, which, while lower than its performance on seen applications, is still remarkably high and significantly surpasses the BC baseline’s 42.1%. This result is central to our paper’s contribution. The BC model tends to overfit to the specific visual and structural patterns present in the training applications. When faced with a new UI, even small deviations can cause it to fail. In contrast, GUI-Learner, having been trained with offline RL on a diverse set of its own experiences, learns a more robust and abstract policy. It learns to focus on the functional goal rather than memorizing specific layouts. For instance, it learns the general concept of ”finding a login button,” regardless of whether the button is blue, green, on the left, or on the right. This ability to acquire generalizable skills is a key objective for foundation models for decision making [1] and is essential for building truly generalist agents that can adapt to the open world [44]. Our results are a promising step in this direction, echoing the goals of comprehensive agent benchmarks like GAIA [40] and AgentBench [39].
@@ V-D Qualitative Analysis and Discussion
Despite its strong performance, GUI-Learner is not infallible. We identified two primary failure modes. The first occurs in tasks requiring deep, external world knowledge. For example, a task like ”Find the population of the city shown in the main image and enter it into the search bar” would fail, as our agent has no mechanism to perform external information retrieval, a capability explored in tool-using agents [27, 50]. The second failure mode involves highly customized or game-like interfaces where standard UI elements are absent. For instance, it struggles to operate a custom video editing timeline that uses non-standard graphical icons for controls. This highlights a limitation of its VLM backbone, which is pre-trained primarily on web and real-world images, not bespoke software UIs. Benchmarks like MathVista [38] and AVIS [30] are pushing the boundaries of what models can understand in such complex visual contexts.
Furthermore, our work shares limitations with the broader field of autonomous agents. The safety and reliability of these agents are paramount. An agent making an error in a banking application could have serious consequences. This points to the need for robust validation, testing frameworks [46], and potentially new security paradigms to prevent malicious attacks, which could even target the underlying physical sensing layers in future multi-modal agents [47]. At the same time, broader analyses of foundation models for decision making highlight efficiency constraints and deployment considerations [1]. The efficiency of the models is also a concern, and while our model is effective, deploying such large models on consumer devices would require significant model compression, an area where techniques like multi-objective convex quantization could be beneficial [51].
@@
Our results demonstrate that the combination of a structured UI representation and a hybrid learning strategy enables the creation of a highly capable and generalizable UI agent. Our work contrasts with approaches that rely on generating code or scripts [52, 54], as our method can operate on any application that presents a visual interface, without needing access to its underlying code.
Future work should proceed in several exciting directions. First, integrating a memory module to handle extremely long-horizon tasks that span minutes or hours is a logical next step. Second, enhancing the agent with tool use capabilities [27] would allow it to overcome its current limitations regarding external knowledge. Third, exploring more diverse training data, including video tutorials of software use [53], could provide a rich, passive source of learning. Finally, the true frontier lies in multi-modal interaction. Future agents could perceive not just screens, but also user affect through vision [60], leverage screen-content QA to better ground on-screen tasks [16], and sense gestures through non-intrusive modalities like WiFi [40, 41], creating a truly context-aware and responsive assistant. This could even extend to health applications, where sensing modalities can monitor a user’s physical state [1, 58]. Building such systems in a privacy-preserving manner, perhaps using advanced federated learning techniques [61, 62], will be a key challenge for the community. Our work, GUI-Learner, provides a robust foundation upon which these future advancements can be built.

### v2

@@ II-E Benchmarks and Evaluation
Rigorous evaluation is the bedrock of scientific progress. The agent community has made significant strides in developing challenging benchmarks and realistic evaluation environments. For web agents, WebArena [zhou2023webarena] provides a realistic and reproducible environment, while the Mind2Web dataset [deng2023mind2web] offers a diverse set of tasks. The evaluation has since expanded to mobile [zhang2024mobile, rawles2023android] and desktop environments [zhang2024desktopenv]. To assess the general capabilities of foundation models as assistants, comprehensive benchmarks like GAIA and AgentBench [liu2023agentbench] have been proposed. Other benchmarks target specific skills, such as instruction-following with vision [biswas2024avis], mathematical reasoning from visual data [lu2024mathvista], or recognizing micro-actions . The insights from these evaluations are crucial; for example, empirical studies have systematically analyzed the failure modes of current agents on real-world tasks [huang2024empirical]. Furthermore, the application of LLM-based agents to software engineering tasks like UI testing is an active area of investigation [araujo2024can], as is the potential for automating UI-to-code generation [daza2023automating]. Our work contributes to this ecosystem by evaluating our proposed agent on a diverse set of tasks and comparing it against relevant state-of-the-art methods, following the principles laid out in these foundational evaluation efforts. We also draw inspiration from work that transfers knowledge from large-scale, unstructured data like internet videos to embodied control [brohan2023rt], a paradigm that holds promise for teaching agents from the vast repository of software tutorials online. Finally, incorporating human-like interaction patterns, such as those informed by gaze data [zheng2024seeact], represents a promising direction for making agents more intuitive and efficient.
@@ III-C Perception Module: Structured UI Representation
The module is built upon a powerful pre-trained Vision-Language Model (VLM), inspired by recent advancements in multimodal understanding [openai2023gpt4v, hong2023cogagent]. The input to the VLM is a prompt that combines the task goal $g$ and the screenshot $I_{t}$. The model is tasked with identifying all interactable elements on the screen. The output is a structured list of UI elements, $U_{t}$, which forms the agent’s symbolic understanding of the current interface state. Each element $e_{i}$ in this list is a tuple containing its essential properties:
@@ III-D Decision Module: Action Generation
The architecture of this module is a Transformer decoder [wei2022chain]. The input to the Transformer is a sequence of tokens constructed by serializing the information from the Perception Module and the task history. Specifically, the structured representation $U_{t}$ is flattened into a sequence of strings, where each element $e_{i}$ is described as, for example, ‘”[ID1] Button ’Login’ at (120, 300, 200, 340)”‘. This sequence is concatenated with the tokenized goal $g$ and a summary of the recent history $h_{t}$.
@@
  • CLICK($\text{id}_{i}$): Clicks the center of the bounding box $B_{i}$ corresponding to element $\text{id}_{i}$.
  • TYPE($\text{id}_{i}$, text): Types the specified ‘text‘ into the element identified by $\text{id}_{i}$, which should typically be a ‘textbox‘. The text to be typed can be generated by the model or extracted from the goal instruction.
@@
  • $\text{Serialize}(\cdot)$ is the function that converts the structured UI data, goal, and history into a flat sequence of tokens for the Transformer.
@@ III-E2 Phase 2: Offline Reinforcement Learning from Self-Exploration
With the labeled offline dataset $\mathcal{D}_{O}$, we use an offline RL algorithm to update the policy. We specifically choose a conservative, actor-critic algorithm like Implicit Q-Learning (IQL) [chen2021decision], which is well-suited for learning from static datasets without querying the environment. IQL avoids explicitly maximizing the Q-function, which can be unstable with out-of-distribution actions, and instead learns by expecting the Q-values for actions in the dataset to be high. It involves learning a state-value function $V_{\psi}$ and a Q-function $Q_{\phi}$ using an expectile regression loss. The policy $\pi_{\theta}$ is then updated to maximize the exponentiated advantage term, pushing the policy towards actions with higher Q-values relative to the average value of the state. The policy update loss is given by:
@@ IV-A1 Datasets and Environments
  • Mind2Web [deng2023mind2web]: This dataset provides a large collection of tasks across various websites, focusing on generalist web agents. We use a subset of its tasks for both training (behavioral cloning) and zero-shot evaluation to test generalization to unseen websites within the same domain.
  • WebArena [zhou2023webarena]: This benchmark provides a realistic and reproducible web environment with tasks that often require complex reasoning and interaction sequences (e.g., online shopping, content management). We use WebArena primarily for evaluation to test the agent’s ability to handle dynamic and complex web applications.
@@
To demonstrate the platform-agnostic nature of our pixel-based approach, we also conduct a qualitative evaluation on tasks from the Android-in-the-Wild dataset [rawles2023android], showcasing the agent’s ability to operate on a mobile GUI without any architectural changes.
@@ IV-A2 Implementation Details
We use a pre-trained Vision-Language Model as the backbone. Specifically, we initialize our model with weights from CogAgent-VQA-hf [hong2023cogagent], which is optimized for GUI understanding and visual question answering. We further fine-tune this model during the behavioral cloning phase to specialize it for our structured UI representation extraction task. The input image resolution is set to 1120x1120 pixels.
@@ V-A Main Quantitative Results
We evaluated GUI-Learner against a suite of strong baselines to demonstrate its effectiveness. The baselines include: (1) Behavioral Cloning (BC), which represents our model trained only with the supervised learning phase (Phase 1); (2) GPT-4V (Zero-shot), where we use the official OpenAI API with a carefully crafted prompt based on the ReAct framework [yao2023react] to directly generate actions from screenshots; and (3) Pix2Act [sharma2020pix2act], a representative earlier vision-based agent that maps pixels directly to actions. The main results, measured in Task Success Rate (SR) and Action Efficiency (AE), are summarized in Table I.
@@
GPT-4V (Zero-shot) [openai2023gpt4v] 41.3 0.68 35.7 0.61
Pix2Act [sharma2020pix2act] 32.5 0.55 28.1 0.52
@@
As shown in Table I, GUI-Learner achieves state-of-the-art performance across all environments and metrics. On the web environments, combining tasks from Mind2Web [deng2023mind2web] and WebArena [zhou2023webarena], our agent achieves a success rate of 82.4%, a significant improvement of 16.6 percentage points over the strong Behavioral Cloning baseline. This underscores the critical role of the offline reinforcement learning phase. While BC provides a strong initial policy, it struggles with the problem of covariate shift; when the agent makes a small mistake and enters a state not seen in the expert data, it often cannot recover. Our RL phase, by learning from a diverse dataset of self-generated trajectories including failures, equips the agent with a more robust policy capable of navigating out of these unfamiliar states. This ability to self-correct is a hallmark of more advanced agentic frameworks like Reflexion [shinn2023reflexion], and our results confirm its importance in the UI domain.
@@
The results on the desktop environments mirror the trends observed on the web. GUI-Learner maintains a substantial performance gap over the baselines, demonstrating that our approach is platform-agnostic. This is a key advantage of our vision-based methodology, which, unlike approaches reliant on web-specific structures like HTML, can seamlessly transition between web, desktop [zhang2024desktopenv], and even mobile environments [zhang2024mobile]. The challenges in desktop automation are distinct, often involving interaction between multiple applications and the operating system shell, a direction explored in recent pilot studies [cheng2024windows]. Our strong performance here suggests that GUI-Learner is a significant step towards a truly universal computer control agent [cradle2024].
@@ V-B Ablation Studies
The most significant finding from our ablation study is the impact of the offline reinforcement learning phase. The ”w/o Offline RL Phase” variant, which is equivalent to our BC baseline, sees its success rate drop by 17.4 points. This confirms our hypothesis that while imitation learning provides an excellent starting point, it is insufficient for achieving high robustness. The RL phase allows the agent to learn from its own mistakes, effectively expanding its knowledge beyond the limited scope of the expert demonstration dataset $\mathcal{D}_{E}$. This self-improvement loop, where an agent learns from its own experience, is a central theme in modern agent research, whether it’s through explicit reflection [shinn2023reflexion] or interaction with a human in the loop [zhang2024self]. Our results provide strong evidence for its efficacy in the context of GUI automation.
@@
In the ”w/o Structured Representation” variant, we replaced our two-module design with a single end-to-end model that maps raw pixels to action coordinates, similar in spirit to older methods [sharma2020pix2act] but built on a modern VLM backbone. The performance plummets to a 45.5% success rate. This drastic drop highlights the critical role of the Perception Module’s abstraction. By converting pixels into a symbolic, object-centric representation, we provide the Decision Module with a much cleaner and more structured input space. This abstraction allows the policy to be invariant to trivial visual changes (e.g., color themes, font styles, minor layout shifts) and to focus on the functional semantics of the interface. This mirrors findings in related domains, where structured representations like graph inference or program synthesis [li2024octopus] have been shown to improve reasoning and generalization.
@@
The ”w/o Expert Pre-training” variant represents an attempt to train the agent using pure reinforcement learning from scratch. The result is a near-complete failure, with a success rate of only 12.8%. This is expected in a domain with such a vast state space and sparse rewards. Without the initial guidance provided by behavioral cloning, the agent’s random exploration is highly unlikely to stumble upon a successful trajectory for any non-trivial task. This finding underscores the sample inefficiency of pure RL in complex interactive domains and validates our choice of a hybrid learning strategy. Bootstrapping from expert data is a common and effective technique, not just in UI automation [li2023web] but also in robotics, where learning from internet-scale video data provides a similar warm start [brohan2023rt].
@@ V-C Generalization to Unseen Applications
GUI-Learner achieves a generalization score of 68.9%, which, while lower than its performance on seen applications, is still remarkably high and significantly surpasses the BC baseline’s 42.1%. This result is central to our paper’s contribution. The BC model tends to overfit to the specific visual and structural patterns present in the training applications. When faced with a new UI, even small deviations can cause it to fail. In contrast, GUI-Learner, having been trained with offline RL on a diverse set of its own experiences, learns a more robust and abstract policy. It learns to focus on the functional goal rather than memorizing specific layouts. For instance, it learns the general concept of ”finding a login button,” regardless of whether the button is blue, green, on the left, or on the right. This ability to acquire generalizable skills is a key objective for foundation models for decision making [fu2024foundation] and is essential for building truly generalist agents that can adapt to the open world [yuan2024towards]. Our results are a promising step in this direction, echoing the goals of comprehensive agent benchmarks like GAIA and AgentBench [liu2023agentbench].
@@ V-D Qualitative Analysis and Discussion
Despite its strong performance, GUI-Learner is not infallible. We identified two primary failure modes. The first occurs in tasks requiring deep, external world knowledge. For example, a task like ”Find the population of the city shown in the main image and enter it into the search bar” would fail, as our agent has no mechanism to perform external information retrieval, a capability explored in tool-using agents [zhou2024ok]. The second failure mode involves highly customized or game-like interfaces where standard UI elements are absent. For instance, it struggles to operate a custom video editing timeline that uses non-standard graphical icons for controls. This highlights a limitation of its VLM backbone, which is pre-trained primarily on web and real-world images, not bespoke software UIs. Benchmarks like MathVista [lu2024mathvista] and AVIS [biswas2024avis] are pushing the boundaries of what models can understand in such complex visual contexts.
Furthermore, our work shares limitations with the broader field of autonomous agents. The safety and reliability of these agents are paramount. An agent making an error in a banking application could have serious consequences. This points to the need for robust validation, testing frameworks [araujo2024can], and potentially new security paradigms to prevent malicious attacks, which could even target the underlying physical sensing layers in future multi-modal agents [10339891]. At the same time, broader analyses of foundation models for decision making highlight efficiency constraints and deployment considerations [fu2024foundation]. The efficiency of the models is also a concern, and while our model is effective, deploying such large models on consumer devices would require significant model compression, an area where techniques like multi-objective convex quantization could be beneficial .
@@
Our results demonstrate that the combination of a structured UI representation and a hybrid learning strategy enables the creation of a highly capable and generalizable UI agent. Our work contrasts with approaches that rely on generating code or scripts [daza2023automating, li2024octopus], as our method can operate on any application that presents a visual interface, without needing access to its underlying code.
Future work should proceed in several exciting directions. First, integrating a memory module to handle extremely long-horizon tasks that span minutes or hours is a logical next step. Second, enhancing the agent with tool use capabilities would allow it to overcome its current limitations regarding external knowledge. Third, exploring more diverse training data, including video tutorials of software use [brohan2023rt], could provide a rich, passive source of learning. Finally, the true frontier lies in multi-modal interaction. Future agents could perceive not just screens, but also user affect through vision [10916982], leverage screen-content QA to better ground on-screen tasks [li2024screenvqa], and sense gestures through non-intrusive modalities like WiFi [9075376], creating a truly context-aware and responsive assistant. This could even extend to health applications, where sensing modalities can monitor a user’s physical state [fu2024foundation, 10699348]. Building such systems in a privacy-preserving manner, perhaps using advanced federated learning techniques [wang2023survey, 10508334], will be a key challenge for the community. Our work, GUI-Learner, provides a robust foundation upon which these future advancements can be built.